(a) cost of goods completed during February (b) cost of the 28 February work in process inventory. Repeat requirement 1 using the FIFO method. 2 Weighted average cost and FIFO: manufacturer Sherida n Textiles Ltd manufactures natural fabrics for the clothing industry. The following data e to the Weaving Department for the month of September: ES27 21 relate ILO3] FIFO 40 Weighted average 60000 52000 Total equivalent units of direct material Total equivalent units of conversion Units completed and transferred out during 50000 50000 September The cost data for September were as follows Work in process 1 September Direct material $94000 44400 during Direct material $164000 272 800 There were 20000 units in process in the Weaving Department on 1 September (100 per cent complete as to direct material and 40 per cent complete as to conversion). Required: Calculate each of the following amounts using weighted average process costing: a) Cost of goods completed and transferred out of the Weaving Department during September. b) Cost of the 30 September work in process inventory in the Weaving Department. Repeat requirement 1 using the FIFO method. 2
Answer #1

Weighted average

Cost per equivalent units
Particulars Material Conversion cost total
As at beginning 94000 44400 138400
Added during the month 164000 272800 436800
Cost to be accounted for 258000 317200 575200
Total eqivalent units 60000 52000
Cost per equivalent unit 4.30 6.10 10.40
Cost per equivalent unit = Cost to be accounted for / Total equivalent units

Cost of goods sold = 50000*10.40 = 520000

Work in process ending inventory = 575200-520000= 55200

FIFO

FIFO method
Beggining units 20000 Transferred out 50,000
Started into production 40000 Ending units 10,000
60000 60000
Equivalent units Material Conversion
Beginning units 20000 20000
Completion in current period 0% 60%
A 0 12000
Units started and completed 30000 30000
Completion 100% 100%
B 30000 30000
Ending Units 10,000 10,000
Completion 100% 20%
C 10000 2000
Total units A+B+C 40000 44000
Cost per equivalent units
Per unit cost Cost Total units Per unit cost
Material cost 164000 40000 4.10
Conversion cost 272800 44000 6.20
Total 436800
Cost of goods sold Material convertion cost
Beginning work in progress A 94000 44400 138400
Beginning inventory completed
Completion 0% 60%
Per unit cost 4.10 6.20
Units 20000 20000
Total cost   B 0 74400 74400
Units started and completed
Completion 100% 100%
Per unit cost 4.10 6.20
Units 30000 30000
Total cost    C 123000 186000 309000
Total cost A+B+C 217000 260400 521800
Ending work in progress
Completion 100% 20%
Per unit cost 4.10 6.20
Units 10,000 10,000
Total cost     41000 12400 53400
Total cost accounted for 575200